### v1

@@ 3.1 Tests and heatmaps
Compared to other products, P3ns was more intense in green/leafy and earthy flavours, more bitter, and more astringent, whereas P4ns had a distinctive tomato ketchup note (Ng et al, 2012). If consumers in the blind sensory evaluation did not expect these sensory properties, it could explain why these two niche products elicited negatively valenced terms more often than other products. Although products did not differ significantly in citation percentages for Angry based on the univariate test results (Test 2; Table 2, Fig. 3), P3nw elicited higher-than-median citation percentages for Angry based on elementwise tests (Test 3; Fig. 4) and univariate pairwise tests (Test 5; results not shown). Again, these results might be explained by the sensory profile of P3nw mentioned above.
@@
Test 4 investigates which products pairs differ considering all $T$ terms simultaneously. For each of the $55$ paired comparisons, the test statistic is the median absolute paired difference across the $T$ terms. Permutation testing (Section 2.3) yields a p-value for each product pair. The critical value that controlled FDR at 0.05 (Section 2.4) was $0.0281$. Based on this critical value, 34 paired comparisons were significantly different. Table 3 shows the multivariate paired comparison test statistics, emphasising these significant differences. The rows and columns of this table are ordered as in Figure 4.
@@
Figure 5 shows the dendrogram from the complete-linkage cluster analysis of products on horizontal and vertical axes of this heatmap, the same dendrogram that appeared in Figure 4 on the left side. The heatmap combined with the dendrogram facilitates the interpretation. It shows that every product cluster 1 member differs significantly from every product cluster 2 member. It also reveals significant differences between some members of product cluster 2, which we describe next.
Figure 5: Heatmap of p-values for the 55 multivariate product paired comparisons in Table 3 (Test 4). With the false discovery rate controlled at 0.05, 34 of the 55 paired were significantly different (shown in green). The diagonal and cells for non-significant tests results are shown in light grey.
Multivariate paired comparison test results showed one of the niche products without added sugar (P3nw) differs significantly from the standard product without added sugar (P2so). Test results also showed P10eo differed significantly from one niche product (P4nw) and two standard products without added sugar (P2so and P8so). Although we can venture reasonable guesses why each of these product pairs differed by referring to the elementwise test results (Figure 4), a better understanding is obtained from univariate product paired comparisons, which will be evaluated next.
@@
The p-values for Happy and Sickly are visualised in heatmaps in Figure 6. In these heatmaps, the direction of the difference in citation percentages is obtained by subtracting the column product from row product, where positive and negative values are colour-coded blue and red, respectively, for product pairs that differ significantly. The differences above and below each main diagonal have the opposite sign but the same absolute magnitude, hence the blue–red symmetry across the diagonal. Products are ordered and shown together with the dendrograms as in Figure 5, which show the results from the complete-linkage cluster analysis of products.
@@
Results show nearly all members of product cluster 1 have significantly higher Happy citation percentages than members of product cluster 2. No products in cluster 2 had a significantly higher citation rate than any product in product cluster 1, but some other within-cluster differences were observed. In product cluster 1, which contains only standard products with added sugar, P7sw was cited more often for Happy than P5sw. In product cluster 2, P2so, which is a standard product without added sugar, was cited more often for Happy than either P3nw, a niche product with added sugar, or P10eo, an economical product without added sugar previously described as catty.
@@
3.2 Dimension reduction using L1-norm PCA
The L1-norm PCA biplot (Figure 7) gives a compact summary of the data structure and shows several results that have been observed in the permutation tests. The scale of these two components is percentage units. The biplot shows scores (products, as points) visualised together with the loading vectors, which are rescaled for legibility. The origin coincides with product P2so, which lies at the median.
Figure 7: L1-norm PCA of the CATA table, with 95% confidence ellipses for the product coordinates.
Figure 8: A. Scree plot of the L1-norm PCA, showing the gains in explaining the CATA table for solutions of increasing dimension. B. Scatterplot of CATA table percentages (horizontal axis) and the values reconstructed from the two-dimensional L1-norm PCA biplot of Figure 7.
Based on Eq. (3), the proportion of the CATA table explained in two components is 0.657. The proportion of the CATA table explained by one L1-norm principal component is 0.586, so the gain in the two-dimensional solution is 0.071. Figure 8A shows the equivalent of a scree plot for the L1-norm PCA, that is, the gains in explanation of the CATA table for increasing dimensions. To show the quality of the L1-norm PCA two-dimensional solution, the concordance between the CATA table percentages $\bf E$ and the estimated values obtained from the two-dimensional L1-norm PCA solution is shown in the scatterplot of Figure 8B. Notice in Eq. (2) that $\bf X$ is median-centered, hence the medians have to be added to ${\bf T}_{K}{\bf P}_{K}^{\sf T}$ to obtain the PCA estimates for the solution in $K=2$ dimensions.
@@ 3.2 Dimension reduction using L1-norm PCA
In most cases, the significant multivariate paired product differences (Figure 5) have non-overlapping ellipses in the L1-norm PCA plane shown (Figure 7). For example, both Test 4 and the L1-norm PCA indicate particular product pairs differ: P10eo and P2so, P3nw and P2so, and P10eo and P8so. In Fig. 7, ellipses for P2so and P8so do not overlap with the ellipse for P10eo (which had catty notes), even though all of these products were in product cluster 2.
The five standard products with added sugar are well separated from the other six products, with the exception of slightly overlapping ellipses for P9sw and P2so, which are better separated in higher dimensions. Table 3, however, confirms the strict separation of these two product clusters statistically.
In some cases, interpretations from the two analyses differ. For example, the multivariate pairwise test (Figure 5) did not detect a significant difference between P1sw and either P5sw or P11sw, even though their 95% confidence ellipses are separated in the first L1-norm PCA plane (Figure 7). The discrepancy probably occurs because the multivariate pairwise test considers all attributes simultaneously, whereas the L1-norm PCA results in this plane considers only the linear combinations of attributes in the two components shown.
@@ 4.2 Comparison with conventional tests
Among the proposed permutation tests in Table 1, the univariate test (Test 2) and the univariate pairwise test (Test 5) are respectively analogous to the Cochran’s Q test (Cochran, 1950; Meyners et al, 2013) and McNemar’s test (McNemar, 1949; Meyners et al, 2013). Results from these two tests are provided in Supplementary Material to facilitate comparisons with permutation testing results in Section 3.
Under a true null hypothesis that $P$ products elicit the same citations proportion, Cochran’s Q test statistic follows the asymptotic chi-squared distribution with $P-1$ degrees of freedom. Cochran’s Q test detected product differences in 30 terms, more than were detected using the permutation testing approach (22). Although controlling FDR will tend to reduce the number of significant test results, in this data set, the number of significant Cochran’s Q test results was the same with and without FDR control, and the number of significant results from Test 2 was also the same with and without FDR control (Table S1).
We also conducted McNemar’s test to determine whether two products elicit the same citation proportion for each of the 55 product pairs on each of the 34 attributes. In each of these 1870 tests, differentiating citations (i.e. in which the term is endorsed for only one of the two products) will be binomially distributed. With FDR controlled at level 0.05, McNemar’s test detected 434 significant paired differences, fewer than were reported based on the permutation testing approach (487). When FDR was uncontrolled, McNemar’s test detected 639 significant paired differences, fewer than were detected using the permutation testing approach (660). These results show that the permutation testing approach is at least sometimes more powerful than the exact version of McNemar’s test. An advantage of McNemar’s test, however, is that it returns consistent output rapidly from an exact distribution that does not require a permutation testing procedure (Table S2).
Previously, Meyners et al (2013) also proposed a global test similar to Test 1 based on the randomisation testing procedure, which was approximated by the sum of Cochran’s Q test statistics. This test is often highly significant in typical consumers tests, as in this data set (Fig. 2).
@@ 4.3 Future studies and limitations
The proposed permutation tests might be applied to other types of data also. For example, quantitative sensory data collected on rating scales could be converted to percentages, then analysed using method proposed in Table 1, where the L1-norm PCA can provide dimension reduction and summary. Other approaches, such as taxicab correspondence analysis (Choulakian, 2006), might also be investigated further.
The method for constructing confidence ellipses in the L1-norm PCA could be investigated further to determine whether these ellipses have approximately frequentist properties and whether other confidence-ellipse constructing procedures are superior (Castura et al, 2022, 2023).
@@
A seed should be set before conducting permutation tests to ensure reproducibility. Software referenced in this manuscript can facilitate such matters. Permutation testing requires enough permutations to achieve stable results. Our results based on 9 999 permutations were usually close (within ±$\pm 0.01$) to results based on 99 999 permutations. Researchers are advised to determine whether results are stable by conducting additional permutation or confirming the conclusions are the same if a different seed is used.
Another potential drawback with the proposed method of clustering terms is that the potential for ties increases when there are relatively few products. Usually, there are more terms than products, but if the number of terms is small, ties might arise when clustering products also.
@@ 5 Conclusion
Our new framework for analysing CATA data is simple and robust. This framework emerges from the principle that every citation in a CATA product-by-term table counts the same as any other citation, no matter the assessor, product, or term. This “one citation, one vote” principle implies a different set of statistical measures based on the L1 norm, contrary to other methods based on the L2 norm. When using the L1 norm, the measure of dispersion is the absolute differences in the CATA values, which are reported in percentage units. In this case, data are analysed “as is”, meaning there are no potentially controversial data transformations. For these reasons, the proposed analyses are often easier to understand than analyses based on the L2 norm where normalisations are often applied to weight terms according to their marginal citation rates and where dispersion based on squared differences gives greater weight to outliers. This approach introduces a considerable simplification into the analysis of CATA tables for the food researcher.
The methods described in this paper are being introduced into the R package cata (Castura, 2025), which will also include this paper’s data on blackcurrant juices.
@@
Chaya, Castura & Greenacre
R code for replicating results
@@
unlist(lapply(results, function(df) {
@@
The L1-norm PCA results were obtained using function l1pca() in the R package pcaL1 (Jot et al, 2023). The following code shows how L1-norm PCA can be run.
@@
squashes.L1pca <- l1pca(results\$CATA.table, projDim = 2)
@@
Confidence ellipses using bootstrap scores can be added using the function CIplot_biv() in the R package easyCODA (Greenacre, 2018).
Cochran’s Q test results
Cochran’s Q test (Cochran, 1950; Meyners et al, 2013) was conducted via the R package cata (Castura, 2025) using the following code:
require(cata)
t(apply(squashes, 3, cochranQ, quiet=TRUE))
@@
According to this table, Cochran’s Q test detected differences among products on 30 of 34 terms. The permutation tests based on Test 2 (Table 2) detected differences among products on 22 of these terms. The eight additional terms for which Cochran’s Q test detected product differences were Desire, Bored, Angry, Resentment, Warm, Reminiscence, Sceptical, and Curious.
McNemar’s test results
McNemar’s test (McNemar, 1949; Meyners et al, 2013) was conducted via the R package cata (Castura, 2025) using the following code:
require(cata)
apply(squashes, 3, mcnemarQ, quiet=TRUE, simplify=FALSE)
McNemar’s test results are presented in Suppl. Table S2. Controlling the FDR at 0.05 yielded a lower BH critical value ($0.0118$) than was obtained in Test 5 ($0.0130$). McNemar’s test also detected fewer significant product pairs ($442$) than those reported for Test 5 ($487$) in Section 3.1.

### v2

@@ 3.1 Tests and heatmaps
Compared to other products, P3nw was more intense in green/leafy and earthy flavours, more bitter, and more astringent (Ng et al., 2012). If consumers in the blind sensory evaluation did not expect these sensory properties, it could explain why this niche product elicited negatively valenced terms more often than other products. Individual tests (Tests 3 and 5) can be significant even when an overall test (Test 2) is not. Some researchers might enforce greater consistency by conducting Test 3 only for terms that were significant based on Test 2, but conducting all tests helps in the interpretation of the consumers’ response elicited by P3nw, adding some cues for the identification of specific differences. Although products did not differ significantly in citation percentages for Angry based on the univariate test results (Test 2; Table 2, Fig. 3), P3nw elicited higher-than-median citation percentages for Angry based on elementwise tests (Test 3; Fig. 4) and univariate pairwise tests (Test 5; results not shown). Again, these results might be explained by the sensory profile of P3nw mentioned above.
@@
Test 4 investigates which product pairs differ considering all $T$ terms simultaneously. For each of the $55$ paired comparisons, the test statistic is the median absolute paired difference across the $T$ terms. Permutation testing (Section 2.3) yields a p-value for each product pair. The critical value that controlled FDR at 0.05 (Section 2.4) was $0.0281$. Based on this critical value, 34 paired comparisons were significantly different. Table 3 shows the multivariate paired comparison test statistics, emphasising these significant differences. The rows and columns of this table are ordered as in Fig. 4.
@@
Fig. 5 shows the dendrogram from the complete-linkage cluster analysis of products on horizontal and vertical axes of this heatmap, the same dendrogram that appeared in Fig. 4 on the left side. The heatmap combined with the dendrogram facilitates the interpretation. It shows standard products with added sugar (product cluster 1) differ significantly from products without added sugar and niche products (product cluster 2). There was no evidence of significant paired differences within product cluster 1, but various paired differences were found within product cluster 2. Among products without added sugar, the economical product P10eo had a different emotional profile than the standard products P8so and P2so. The two niche products, P3nw and P4nw, elicited a different emotional response than products P2so and P10eo, respectively, from the same product cluster 2.
Figure 5: Heatmap of p-values for the 55 multivariate product paired comparisons in Table 3 (Test 4). With the false discovery rate controlled at 0.05, 34 of the 55 pairs were significantly different (shown in green). The diagonal and cells for non-significant tests results are shown in light grey.
Test results also showed P10eo differed significantly from one niche product (P4nw) and two standard products without added sugar (P2so and P8so). Although we can venture reasonable guesses why each of these product pairs differed by referring to the elementwise test results (Fig. 4), a better understanding is obtained from univariate product paired comparisons, which will be evaluated next.
@@
The p-values for Happy and Sickly are visualised in heatmaps in Fig. 6. In these heatmaps, the direction of the difference in citation percentages is obtained by subtracting the column product from the row product (i.e. rows minus columns), where positive and negative values are colour-coded blue and red, respectively, for product pairs that differ significantly. The differences above and below each main diagonal have the opposite sign but the same absolute magnitude, hence the blue–red symmetry across the diagonal. Products are ordered and shown together with the dendrograms as in Fig. 5, which show the results from the complete-linkage cluster analysis of products.
@@
Results show nearly all members of product cluster 1 have significantly higher Happy citation percentages than members of product cluster 2. No products in cluster 2 had a significantly higher citation rate than any product in product cluster 1, but some other within-cluster differences were observed. The observed Happy citation percentages were larger in product cluster 1 than in product cluster 2. Within product cluster 1, the observed Happy citation percentages were highest for P7sw and lowest for P5sw, but these differences were not significant. Due to these differences, we detected more (6) significant paired differences between P7sw and the products in product cluster 2 than detected (3) between P5sw and products in product cluster 2. P3nw and P10eo had the lowest Happy citation percentages (both 15%); both these products differed significantly from all five products in product cluster 1.
@@
3.2 Dimension reduction using L1-PCA
The L1-PCA biplot (Fig. 7) gives a compact summary of the data structure and shows several results that have been observed in the permutation tests. The scale of these two components is percentage units. The biplot shows scores (products, as points) visualised together with the loading vectors, which are rescaled for legibility. The origin coincides with product P2so, which lies at the median.
Figure 7: L1-PCA biplot of the CATA table, with 95% confidence ellipses for the product medians.
Figure 8: A. Scree plot of the L1-PCA, showing the gains in explaining the CATA table for solutions of increasing dimension ($K$). B. Scatterplot of CATA table percentages (horizontal axis) and the values reconstructed from the two-dimensional L1-PCA biplot of Fig. 7.
Based on Eq. (3), the proportion of the CATA table explained in two components is 0.657. The proportion of the CATA table explained by one L1-PCA component is 0.586, so the gain in the two-dimensional solution is 0.071. Fig. 8A shows the equivalent of a scree plot for the L1-PCA, that is, the gains in explanation of the CATA table for increasing dimensions. To show the quality of the L1-PCA two-dimensional solution, the concordance between the CATA table percentages $\bf E$ and the estimated values obtained from the two-dimensional L1-PCA solution is shown in the scatterplot of Fig. 8B. Notice in Eq. (2) that $\bf X$ is median-centered, hence the medians have to be added to ${\bf T}_{K}{\bf P}_{K}^{\sf T}$ to obtain the PCA estimates for the solution in $K=2$ dimensions.
@@ 3.2 Dimension reduction using L1-PCA
In most cases, the significant multivariate paired product differences (Fig. 5) have non-overlapping ellipses in the L1-PCA plane shown (Fig. 7). For example, both Test 4 and the L1-PCA indicate particular product pairs differ: P10eo and P2so, P3nw and P2so, and P10eo and P8so. In Fig. 7, ellipses for P2so and P8so do not overlap with the ellipse for P10eo (which had catty notes), even though all of these products were in product cluster 2.
The five standard products with added sugar are well separated from the other six products, with the exception of slightly overlapping ellipses for P9sw and P2so, and P9sw and P4nw, which are better separated in higher dimensions. Table 3, however, confirms the strict separation of these two product clusters statistically.
In some cases, interpretations from the two analyses differ. For example, the multivariate pairwise test (Fig. 5) did not detect a significant difference between P1sw and either P5sw or P11sw, even though their 95% confidence ellipses are separated in the first L1-PCA plane (Fig. 7). The discrepancy probably occurs because the multivariate pairwise test considers all attributes simultaneously, whereas the L1-PCA results in this plane considers only the linear combinations of attributes in the two components shown.
@@ 4.2 Comparison with conventional tests
Among the proposed permutation tests in Table 1, the univariate test (Test 2) and the univariate pairwise test (Test 5) are respectively analogous to the Cochran’s Q test (Cochran, 1950; Meyners et al., 2013) and McNemar’s test (McNemar, 1949; Meyners et al., 2013). Results from these two tests are provided in Supplementary Material to facilitate comparisons with permutation testing results in Section 3.
Under a true null hypothesis that $P$ products elicit the same citations proportion, Cochran’s Q test statistic follows the asymptotic chi-squared distribution with $P-1$ degrees of freedom. Cochran’s Q test detected product differences in 30 terms, more than were detected using the permutation testing approach (22). Although controlling FDR will tend to reduce the number of significant test results, in this data set, the number of significant Cochran’s Q test results was the same with and without FDR control, and the number of significant results from Test 2 was also the same with and without FDR control (Supplementary Material, Table S1).
We also conducted McNemar’s test to determine whether two products elicit the same citation proportion for each of the 55 product pairs on each of the 34 attributes. In each of these 1870 tests, differentiating citations (i.e. in which the term is endorsed for only one of the two products) will be binomially distributed. With FDR controlled at level 0.05, McNemar’s test detected 434 significant paired differences, fewer than were reported based on the permutation testing approach (487). When FDR was uncontrolled, McNemar’s test detected 639 significant paired differences, fewer than were detected using the permutation testing approach (660). These results show that the permutation testing approach is at least sometimes more powerful than the exact version of McNemar’s test. An advantage of McNemar’s test, however, is that it returns consistent output rapidly from an exact distribution that does not require a permutation testing procedure (Supplementary Material, Table S2).
Previously, Meyners et al. (2013) also proposed a global test similar to Test 1 based on the randomisation testing procedure, which was approximated by the sum of Cochran’s Q test statistics. This test is often highly significant in typical consumers tests, as in this data set (Fig. 2).
@@ 4.3 Future studies and limitations
The proposed permutation tests might be applied to other types of data also. For example, quantitative sensory data collected on rating scales could be converted to percentages, then analysed using method proposed in Table 1, where the L1-PCA can provide dimension reduction and summary. Other approaches, such as taxicab correspondence analysis (Choulakian, 2006), might also be investigated further.
The method for constructing confidence ellipses in the L1-PCA could be investigated further to determine whether these ellipses have approximately frequentist properties and whether other confidence-ellipse constructing procedures are superior (Castura et al., 2022, 2023).
@@
A seed should be set before conducting permutation tests to ensure reproducibility. Software referenced in this manuscript can facilitate such matters. Permutation testing requires enough permutations to achieve stable results. Our results based on 9 999 permutations were usually close (within ±$\pm 0.01$) to results based on 99 999 permutations. Researchers are advised to determine whether results are stable by conducting additional permutations or confirming that the conclusions are the same if a different seed is used.
Neither term clusters nor product clusters seemed to be influenced by marginal citation percentages of the terms and products, respectively; however, a potential drawback with the proposed method of clustering terms is that ties might increase when there are relatively few products. Usually, there are more terms than products, but if the number of terms is small, ties might arise when clustering products also.
@@ 5 Conclusion
Our new framework for analysing CATA data is simple and robust. This framework emerges from the principle that every single citation recorded in a CATA product-by-term table counts the same as any other citation, no matter the assessor, product, or term. This “one citation, one vote” principle implies a different set of statistical measures based on the L1 norm, contrary to other methods based on the L2 norm. When using the L1 norm, the measure of dispersion is based on the absolute differences in the CATA values, which are reported in percentage units, whereas methods based on the L2 norm are based on squared differences, often with some type of normalization. In the present approach, data are analysed “as is”, meaning there are no potentially controversial data transformations. Thus, even though there are differences in the frequencies of checking between terms and between assessors, we avoid any attempt to normalize the data, both term-wise and assessor-wise, since such normalizations would imply that citations get varying numbers of “votes”, depending on the terms and/or the assessor.
Thanks to the absence of any transformation on the CATA table, the proposed analyses are often easier to understand than analyses based on the L2 norm. Our approach using the L1 norm introduces a considerable simplification into the analysis of CATA tables for the food researcher.
The methods described in this paper are being introduced into the R package cata (Castura, 2025), which will also include this paper’s data on blackcurrant squashes.
@@
Supplementary material is provided in the form of two files.
  1. 1.
    A file with five sections: (S1) R code for replicating results, (S2) Cochran’s Q test results, (S3) McNemar’s test results, (S4) principal component analysis and correspondence analysis, and (S5) a description of the 3D video animation of the L1-PCA results in Suppl. Video V1.
  2. 2.
    Suppl. Video V1. Score plot for L1-PCA of the CATA table with three components showing 95% confidence ellipoids for the product coordinates. The video shows rotation around the vertical second axis, so the dispersion of the product ellipsoids on the third dimension can be seen.
Supplementary Material
@@
A new approach for analysing check-all-that-apply (CATA) data
using L1-norm methods
Carolina Chaya, John C. Castura and Michael J. Greenacre
S1. R code for replicating results
@@
unlist(lapply(res, function(df) {
@@
The L1-PCA results were obtained using the R package pcaL1 (Jot et al., 2023). The following code shows how L1-PCA can be run.
@@
squashes.L1pca <- l1pca(res\$CATA.table, projDim = 2)
@@
After obtaining bootstrap scores, confidence ellipses can be added using the R package easyCODA (Greenacre, 2018).
S2. Cochran’s Q test results
@@
In total, Cochran’s Q test (Suppl. Table S1) detected differences among products on 30 of 34 terms. The permutation tests based on Test 2 (Table 2) detected differences among products on 22 of these terms. The eight additional terms for which Cochran’s Q test detected product differences were Desire, Bored, Angry, Resentment, Warm, Reminiscence, Sceptical, and Curious.
S3. McNemar’s test results
McNemar’s test results are presented in Suppl. Table S2. Controlling the FDR at 0.05 yielded a lower BH critical value ($0.0118$) than was obtained in Test 5 ($0.0130$). McNemar’s test also detected fewer significant product pairs ($442$) than was reported for Test 5 ($487$) in Section 3.1.
@@
S4. Principal component analysis and correspondence analysis
In this section, we will compare the L1-PCA used in this paper with the conventional PCA which uses the L2 norm, referred to as L2-PCA. We will also discuss different ways of conducting correspondence analysis of a CATA frequency table. Results based on the CATA table of blackcurrant squashes will illustrate the differences between methods.
In the L1-PCA, used in the “one citation, one vote” approach, the deviation of each cell $e_{pt}$ of the CATA table E is measured with respect to the median M($t$) by the absolute difference $|e_{pt}-\textsf{M}(t)|$ (see Eq. (1) of the paper). In L2-PCA, the deviations are measured by squared differences with respect to the term mean $\bar{e}_{t}$: $(e_{pt}-\bar{e}_{t})^{2}$, not with respect to the median. Writing this squared difference as $|e_{pt}-\bar{e}_{t}|\cdot|e_{pt}-\bar{e}_{t}|$, shows that the absolute difference with respect to the term mean, $|e_{pt}-\bar{e}_{t}|$, can be considered weighted by exactly the same amount. This weighting increases the contribution of large deviations from the mean compared to small deviations in the measure of total dispersion in the data table, whereas our L1-norm approach leaves such deviations (from the median) unweighted. Nevertheless, the two-dimensional L2-PCA solution of the blackcurrant squashes table, shown in Fig. S1, is quite similar to that of the L1-PCA result of Fig. 7. Apart from slightly different emphases on some terms, the most noticeable difference is the increased elongation of the confidence ellipses along the first dimension, which accounts for a high percentage (82.3%) of the variance in the table of citation percentages. This high percentage compared to the lower percentage of 58.6% of the first dimension in the L1-PCA is a result of the squaring of the deviations, a convex function that exaggerates the higher dispersions. Notice too that in the L2-PCA, as in the L1-norm version, the term citation values are not standardized, since the natural variances of the terms are required to be preserved, not equalized.
Figure S1: L2-PCA of the CATA table, with confidence ellipses for the product positions. Percentages of explained variance on these two dimensions are 82.3% and 6.6% respectively.
Correspondence analysis (CA) is a popular dimension-reduction method for tables of counts, and inherent in CA is the chi-square normalization. This normalization divides each squared difference between the observed count $e_{pt}$ and respective term mean count $\bar{e}_{t}$ by the inverse of the mean count: $(e_{pt}-\bar{e}_{t})^{2}/\bar{e}_{t}$, which is the typical chi-square form of (observed-expected)2/expected. This quantity can be written as $|e_{pt}-\bar{e}_{t}|\cdot|e_{pt}-\bar{e}_{t}|/\bar{e}_{t}$. So CA, being an L2-norm method, not only “boosts” each absolute difference with respect to the mean by the same absolute difference, but then additionally weights by the inverse of the mean, which has an effect of downweighting the higher counts. Looked at it this way, CA can be thought of as having weights on each deviation $|e_{pt}-\bar{e}_{t}|$ equal to $|e_{pt}-\bar{e}_{t}|/\bar{e}_{t}$, where the numerator and denominator in this weight ratio can sometimes partially cancel out, since terms with larger means will tend to have larger deviations from the mean. Nevertheless, it is clear that this weighting (or any other weighting of the terms) introduces varying numbers of “votes” for each citation, which disagrees with our “one citation, one vote” principle.
However, there is a problem applying regular CA to a CATA table, which is not a contingency table. Using CA in this case has to be done with careful consideration of a major property inherent in the method, namely the expressing of the row (or column) counts relative to their respective totals. For example, the set of row (product) counts is converted to a set of relative counts, called profiles or compositions, and it is these relative counts that are being visualized by CA. These relative counts depend on the total number of citations for each product, which vary. Mahieu et al. (2021) have quite correctly drawn attention to this point and proposed an alternative analysis, multiple response CA (mrCA), which does not change the counts into relative counts, but still conserves the chi-square style of standardization. It turns out that their mrCA can be equivalently computed using either a variant of CA called “subset CA” (Greenacre & Pardo, 2006) or another version called “CA of raw data” (Greenacre, 2010).
Subset CA analyses a selected set of rows and/or columns in a table but maintains the margins of the original table. To keep the original CATA table values and not relativize them with respect to their row totals, a dummy column is added to the CATA table to force the table to have equal row totals. For example, the dummy column could be, for each row $p$, $\sum_{t}(100-e_{pt})$, i.e. the sum of the percentages of non-citations in the CATA table. Then the row sums of the new $P\times(T+1)$ table are a constant $100T$. The subset CA is performed indicating that the subset of columns of interest consists of the first $T$ columns, omitting the dummy column introduced to force equal row margins.
The CA of raw data, which can be called rawCA, originated independently in the context of ecological abundance data, where the actual abundance levels were of interest, not their relative abundances that regular CA analyses by default. Thus, in this analysis the inherent proportioning out of the counts with respect to the totals is suppressed, and the raw data are analysed, not their relative values.
It turns out that the solutions of both mrCA and rawCA are equivalent to the subset CA solution, which was also developed in a completely different context, namely for analysing survey data where specific response options were required to be visualized.
The subset CA analysis can be simply achieved with the R package ca (Nenadic & Greenacre, 2007), using the option subsetcol, as follows:
# Subset CA to perform Mahieu et al.’s CATA analysis
# Suppose CATA table of percentages is in E, with T columns
require(ca)
E2 <- cbind(E, 100*T - rowSums(E))
E.subCA <- ca(E2, subsetcol=1:T)
plot(E.subCA)
The result of the above code, applied to the CATA table, and enhanced with the confidence ellipses, is given in Fig. S2. Again, there are only minor differences compared to the L1-PCA biplot (Fig. 7) and the L2-PCA biplot (Fig. S1). However, these analyses might not be similar in other applications.
Figure S2: Subset CA of the CATA table, with confidence ellipses for the product positions. Percentages of explained variance on these two dimensions are 78.3% and 8.2% resoectively.
S5. Three-dimensional view of the L1-PCA solution
Suppl. Video V1 shows an animation of the three-dimensional L1-PCA solution, rotating by 360 degrees around the vertical second axis. The ellipsoids of the products that are standard with added sugar are coloured red, to show the contrast with the other products on the first dimension. The video pauses at the initial dimension 1 by dimension 2 view and then again at the dimension 3 by dimension 2 view. In the first pausing, the non-overlapping confidence ellipsoids show the products are well separated, but in the second pausing, all confidence ellipsoids overlap on the third dimension, showing that the products are not well distinguished in the third component. We focus our interpretation on the two-component L1-PCA, which differs slightly from the first plane of the three-component L1-PCA solution due to the non-nestedness of L1-PCA.